A missile is shot straight up into the air. At the peak of its trajectory, it breaks up into three pieces of equal mass, all of which move horizontally away from the point of the explosion. One piece travels in a horizontal direction of 35° east of north with a speed of 31 m/s. The second piece travels in a horizontal direction of 11° south of west with a speed of 7 m/s. What is the velocity of the remaining piece?
Answer #1

hsing conservation of momentum

p1 + p2 + p3 = 0

m v1 + m v2 + m v3 = 0

31 ( sin 35 i + cos 35 j) + 7 ( - cos 11i - sin 11j) + v3 = 0

v3 = - 10.91 i - 24.06 j

net magnitude of velocity

v = 26.416 m/s

========